Note to Jesse: Ho is set up, Hb is not observing next experiment (contradictory to calendar)

Hobart 12m

Module: UVLBI-06/ Start Position: 0.224 GBs

  • 17:00 - Experiment started OK (Tiege).
  • 01:00 - rfpic error message switched off. (Lucia)
  • 01:30 - we entered the antenna to see the cable wrap. possible brief power interruption, though unlikely. (Lucia)

Katherine 12m

Module: HOB+0047/ Start Position: 0.256 GBs

  • 17:00 - Experiment started OK (Tiege).
  • 07:15 UT - Maser status unknown, maser computer is no longer connecting to the network (known issue), possibly (according to Jamie) connecting intermittently. Have not been able to connect my entire shift. (Ellen)

Yarragadee 12m

Module: MET-0001/ Start Position: 0.417 GBs

  • 17:00 - Experiment started OK (Tiege).
  • 07:15 UT - S-band (channels 09-14) have been high for most of the time (until now at 13:13UT) but occasionally return to normalish levels. No storm or any other obvious explanation. (Ellen)